# Pricing Online LLM Services with Data-Calibrated Stackelberg Routing Game

### v1

@@ Introduction
(1) Problem formulation (Section Stackelberg Routing Game). We formulate the service pricing problem as a Stackelberg game, where providers act as leaders and customers as followers. Given the initial pricing strategy profile $\mathcal{P}$, service selection strategy profile $\mathcal{F}$, user-defined routing functions, and a target provider $s$, it alternates between updating $\mathcal{F}$ to reach a Nash equilibrium under $\mathcal{P}$, and adjusting $s$’s pricing strategy to maximize its profit. This yields a mathematical program with equilibrium constraints (MPEC), which is usually NP-hard. We prove the existence and uniqueness of the Nash equilibrium.
(2) Game calibration (Section Data-Driven Game Calibration). We calibrate our Stackelberg routing model to real LLM routing data by learning user‑preference parameters $\bm{\theta}$. At any given price and market condition, the user side admits a single Nash equilibrium flow, delivered by a predictor function $\mathcal{F}^{*}(\cdot)$. Fitting the game model to observed traffic allows us to recover latent preferences and align the model with real behaviors.
(3) Game abstraction (Section Dynamic Pricing with Game Abstraction). To reduce complexity in large-scale markets, we aggregate both users and service providers into abstracted groups. On the user side, we group users by the apps they use, based on the assumption that users of the same app share similar preferences. On the provider side, we propose a deep aggregation network that learns to abstract service providers. To prevent overfitting, $\mathsf{PriLLM}$ samples multiple pricing strategies for $s$ and minimizes the discrepancy in selection strategies across them.
(4) Experimental study (Section Evaluation). Using real-life market data from OpenRouter, we empirically find the following. (a) $\mathsf{PriLLM}$ accurately captures real-world market dynamics parameters, achieving a high $R^{2}$ score of 0.8982 in fitting user traffic flows. (b) $\mathsf{PriLLM}$ outperforms pMPEC and Smooth, the best baselines, by 9.2% and 14.3% on average, respectively, up to 11.4% and 17.3%. (c) $\mathsf{PriLLM}$ achieves over 95% of the optimal profit, requiring less than 5% of the computation time of the brute-force optimal solver.
Related Work
Stackelberg Game in Service Pricing
@@
Model Relaxation in MPEC
@@
Stackelberg Routing Game
@@
Data-Driven Game Calibration
@@
Learnability of Game Parameters
@@
The potential function is the sum of these two parts: $\Phi({\mathcal{F}})=\Phi_{\text{Fixed}}({\mathcal{F}})+\Phi_{\text{Congestion}}({\mathcal{F}})$, where $\mathcal{F}$$^{*}$ is the uniqueness solution to minimization of $\Phi({\mathcal{F}};\bm{\theta},\bm{O},\{D_{i}\}_{i\in U})$. A proof is provided in the Section Users’ NE Calculation. To learn $\bm{\theta}$ via gradient-based methods, we also need compute the gradient of a loss function through ${\mathcal{F}}^{*}(\cdot)$.
@@
Initialize and Learn Game Parameters
The learning process is sensitive to the initial values of $\bm{\theta}$, as poor initialization often leads to a failure to converge. Hence, we propose an initialization strategy to get the high quality initial values $\{b_{j}\}$ for $\bm{\theta}$ . The method takes representative days of real-world traffic data (${\mathcal{F}}^{\text{real}}_{t}$) as input, and outputs a initial parameters $\bm{\theta}_{\text{init}}$ , detailed in Section Users’ NE Calculation.
@@
Dynamic Pricing with Game Abstraction
@@
Rival Ranking and Abstraction
@@
Loss Function and Model Solving
@@
where $\phi$ represents the parameters of our deep aggregation network. $X$ is the total number of candidate price points sampled for the target provider $s$. $\mathbf{Y}=[Y_{1},\dots,Y_{L}]$ is the profit vector in which each $Y_{k}$ is the profit for provider $s$ at the $k$-th candidate price, computed in the original routing game. But $\hat{\mathbf{Y}}(\phi)=[\hat{Y}_{1}(\phi),\dots,\hat{Y}_{L}(\phi)]$ is computed in the simplified game generated by $\mathsf{PriLLM}$’s aggregator. $\mathbf{Y}||_{\infty}$ is the L-infinity norm (i.e., the maximum absolute value) of the original profit vector. We use it to normalize both curves, which stabilizes the training process against varying scales of profit. Detail calculation of NE is provided in Section Users’ NE Calculation.
@@
Overall Performance of $\mathsf{PriLLM}$
@@ Acknowledgments
Data Acquisition and Parameter Calibration
@@
User-Side Equilibrium Existence and Uniqueness
We provide detailed proof for Theorem 1 in Section Game Formulation.
@@
Users’ NE Calculation
We provide calculation method of user-side NE for Section Learnability of Game Parameters.
@@
Learnability of Game Parameters
We provide detailed proof for Theorem 2 in Section Learnability of Game Parameters.
@@
Initialization of Subjective Preferences $\{b_{j}\}$
We provide the initialization method of subjective preferences $\{b_{j}\}$ in Section Initialize and Learn Game Parameters.
@@
As formulated, Problem Initialization of Subjective Preferences $\{b_{j}\}$ is a standard LP that can be solved efficiently to obtain the optimal biases $\{b_{j}^{*}\}$. The initial parameter vector for the calibration is therefore set to $\bm{\theta}_{\text{init}}=\{w_{p}=1.0,w_{q}=1.0,w_{d}=1.0,\{b_{j}^{*}\}\}$.

### v2

@@ 1 Introduction
(1) Problem formulation (Section 3). We formulate the service pricing problem as a Stackelberg game, where providers act as leaders and customers as followers. Given the initial pricing strategy profile $\mathcal{P}$, service selection strategy profile $\mathcal{F}$, user-defined routing functions, and a target provider $s$, it alternates between updating $\mathcal{F}$ to reach a Nash equilibrium under $\mathcal{P}$, and adjusting $s$’s pricing strategy to maximize its profit. This yields a mathematical program with equilibrium constraints (MPEC), which is usually NP-hard. We prove the existence and uniqueness of the Nash equilibrium.
(2) Game calibration (Section 4). We calibrate our Stackelberg routing model to real LLM routing data by learning user‑preference parameters $\bm{\theta}$. At any given price and market condition, the user side admits a single Nash equilibrium flow, delivered by a predictor function $\mathcal{F}^{*}(\cdot)$. Fitting the game model to observed traffic allows us to recover latent preferences and align the model with real behaviors.
(3) Game abstraction (Section 5). To reduce complexity in large-scale markets, we aggregate both users and service providers into abstracted groups. On the user side, we group users by the apps they use, based on the assumption that users of the same app share similar preferences. On the provider side, we propose a deep aggregation network that learns to abstract service providers. To prevent overfitting, $\mathsf{PriLLM}$ samples multiple pricing strategies for $s$ and minimizes the discrepancy in selection strategies across them.
(4) Experimental study (Section 6). Using real-life market data from OpenRouter, we empirically find the following. (a) $\mathsf{PriLLM}$ accurately captures real-world market dynamics parameters, achieving a high $R^{2}$ score of 0.8982 in fitting user traffic flows. (b) $\mathsf{PriLLM}$ outperforms pMPEC and Smooth, the best baselines, by 9.2% and 14.3% on average, respectively, up to 11.4% and 17.3%. (c) $\mathsf{PriLLM}$ achieves over 95% of the optimal profit, requiring less than 5% of the computation time of the brute-force optimal solver.
2 Related Work
2.1 Stackelberg Game in Service Pricing
@@ 2.2 LLM Service Evaluation
2.3 Model Relaxation in MPEC
@@
3 Stackelberg Routing Game
@@ 3.2 Game Formulation
4 Data-Driven Game Calibration
@@
4.1 Learnability of Game Parameters
@@
The potential function is the sum of these two parts: $\Phi({\mathcal{F}})=\Phi_{\text{Fixed}}({\mathcal{F}})+\Phi_{\text{Congestion}}({\mathcal{F}})$, where $\mathcal{F}$$^{*}$ is the uniqueness solution to minimization of $\Phi({\mathcal{F}};\bm{\theta},\bm{O},\{D_{i}\}_{i\in U})$. A proof is provided in the Section 9.3. To learn $\bm{\theta}$ via gradient-based methods, we also need compute the gradient of a loss function through ${\mathcal{F}}^{*}(\cdot)$.
@@
4.2 Initialize and Learn Game Parameters
The learning process is sensitive to the initial values of $\bm{\theta}$, as poor initialization often leads to a failure to converge. Hence, we propose an initialization strategy to get the high quality initial values $\{b_{j}\}$ for $\bm{\theta}$ . The method takes representative days of real-world traffic data (${\mathcal{F}}^{\text{real}}_{t}$) as input, and outputs a initial parameters $\bm{\theta}_{\text{init}}$ , detailed in Section 9.3.
@@ 4.2 Initialize and Learn Game Parameters
5 Dynamic Pricing with Game Abstraction
@@
5.1 Rival Ranking and Abstraction
@@
5.2 Loss Function and Model Solving
@@
where $\phi$ represents the parameters of our deep aggregation network. $X$ is the total number of candidate price points sampled for the target provider $s$. $\mathbf{Y}=[Y_{1},\dots,Y_{L}]$ is the profit vector in which each $Y_{k}$ is the profit for provider $s$ at the $k$-th candidate price, computed in the original routing game. But $\hat{\mathbf{Y}}(\phi)=[\hat{Y}_{1}(\phi),\dots,\hat{Y}_{L}(\phi)]$ is computed in the simplified game generated by $\mathsf{PriLLM}$’s aggregator. $\mathbf{Y}||_{\infty}$ is the L-infinity norm (i.e., the maximum absolute value) of the original profit vector. We use it to normalize both curves, which stabilizes the training process against varying scales of profit. Detail calculation of NE is provided in Section 9.3.
@@ 6 Evaluation
6.1 Overall Performance of $\mathsf{PriLLM}$
@@ 9 Appendix
9.1 Data Acquisition and Parameter Calibration
@@
9.2 User-Side Equilibrium Existence and Uniqueness
We provide detailed proof for Theorem 1 in Section 3.2.
@@ 9.2 User-Side Equilibrium Existence and Uniqueness
9.3 Users’ NE Calculation
We provide calculation method of user-side NE for Section 4.1.
@@ 9.3 Users’ NE Calculation
9.4 Learnability of Game Parameters
We provide detailed proof for Theorem 2 in Section 4.1.
@@ 9.4 Learnability of Game Parameters
9.5 Initialization of Subjective Preferences $\{b_{j}\}$
We provide the initialization method of subjective preferences $\{b_{j}\}$ in Section 4.2.
@@ 9.5 Initialization of Subjective Preferences $\{b_{j}\}$
As formulated, Problem 9.5 is a standard LP that can be solved efficiently to obtain the optimal biases $\{b_{j}^{*}\}$. The initial parameter vector for the calibration is therefore set to $\bm{\theta}_{\text{init}}=\{w_{p}=1.0,w_{q}=1.0,w_{d}=1.0,\{b_{j}^{*}\}\}$.